If the intreset value of saving $ 24000 is $ 6000 for a period of 5 years, then find the rate of interest per year.
Answer by jim_thompson5910(35256)

I = Prt

6000 = 24000*r*5

6000 = 120000r

6000/120000 = r

r = 6000/120000

r = 6/120

r = 1/20

r = 0.05

So the interest rate is 5%